# Adversaries With Incentives: A Strategic Alternative to Adversarial Robustness

### v1

@@ Abstract
Adversarial training aims to defend against adversaries: malicious opponents whose sole aim is to harm predictive performance in any way possible—a rather harsh perspective, which we assert results in unnecessarily conservative models. Instead, we propose to model opponents as simply pursuing their own goals, rather than working directly against the classifier. Employing tools from strategic modeling, our approach uses knowledge or beliefs regarding the opponent’s possible incentives as inductive bias for learning. Our method of strategic training is designed to defend against opponents within an ‘incentive uncertainty set’: this resorts to adversarial learning when the set is maximal, but offers potential gains when it can be appropriately reduced. We conduct a series of experiments that show how even mild knowledge regarding the adversary’s incentives can be useful, and that the degree of potential gains depends on how incentives relate to the structure of the learning task.
@@ 1 Introduction
The goal of adversarial learning is to train classifiers that are robust to adversarial attacks, defined as small input perturbations crafted to fool the classifier into making erroneous predictions [41, 15]. As its name suggests, adversarial learning seeks to defend against an adversary—an ominous opponent whose actions are intended to induce maximal harm to the classifier’s performance. Learning in this setting can therefore be modeled as a zero-sum game between the learner and the adversary, in which each player’s gain translates to an equivalent loss for the other. This lends to formulating adversarial training as a minimax optimization problem in which the learner (min) controls model parameters, and the adversary (max) controls bounded-norm additive noise terms that are applied to inputs [28].
@@
Since learning seeks to make correct predictions, adversaries are modeled as gaining from predictive errors of any type. This gives an adversary much flexibility: while there is typically only a single (or perhaps a few) correct labels, most labels are wrong. Thus, to obtain robustness, learning must defend simultaneously against all attacks targeting any incorrect label. And while this approach can in principle provide a high degree of protection, it does so by placing demanding restrictions on the learning objective, whether explicitly or implicitly [37]. This makes robustness attainable—but at a cost; for example, it is well-known that adversarially trained models suffer from deteriorated generalization [38, 45] and reduced performance on clean (i.e., non-adversarial) data [42].
@@
In this paper, we argue that such costs can potentially be reduced by injecting into the learning objective a notion of the opponent’s incentives, i.e., what the opponent wants—which we propose to use as inductive bias. Notice that while conventional adversarial training considers the prospective harm of an attack, it lacks to consider the possible motives behind it: the fact that a panda can be turned into a gibbon does not provide grounds for why (nor whether) an opponent would wish to do so. Alternatively, our working assumption is that realistic opponents are much more likely to simply promote their own self-interests, rather than to always and purposefully counter those of learning—as is assumed in adversarial training. Thus, if we have some beliefs or knowledge regarding the opponent’s interests, even if quite general, then this can and should be exploited to improve prediction.
As an example, consider a malicious opponent who wishes to disrupt traffic by manipulating road signs. Adversarial training, which assumes that attack targets can be arbitrary, will invest effort in protecting against all possible transformations. But note there is structure to the opponent’s incentives: for example, the opponent has more to gain from mapping a ‘speed limit 90’ to a stop sign than to a similar speed limit; should prefer ‘no motor vehicles’ over ‘no motorcycles’ since it is more restrictive (though visually similar); will benefit from ‘no U-turn’ only if a left turn is possible; and should be indifferent at a T-intersection between ‘no turn left’ and ‘must turn right’ since they entail the same outcomes. If such preferences are encoded into the learning objective, then training can focus on protecting against viable attacks, while avoiding the unnecessary costs of defending against irrelevant ones.
Borrowing from the literature on strategic classification [17, 6, 26, 27], our approach is to model the opponent as a strategic agent whose actions are intended to maximize its own utility (rather than to generally minimize accuracy), this lending to a non-zero sum game formulation. We focus on utilities that depend on prediction outcomes, meaning that the opponents we consider have preferences regarding which predictive errors are made. Given an uncertainty set of possible utilities for the opponent (or a distribution over such) which are believed to be plausible for the task at hand, our goal will be to defend against the worst-case opponent in this set, a notion which we refer to as strategic robustness.
@@
Figure 1: Strategic opponents benefit only from certain types of attacks, such as preserving label categories (semantic), changing labels drastically (anti-semantic), or mapping each class to a specific target (1-hot). As such, they offer a flexible middle ground between fully adversarial and clean settings. Our strategic training approach enables to encode prior knowledge regarding the opponent’s possible incentives into the learning objective, this ensuring that training defends only against viable targets.
Our first observation is that standard adversarial training conforms to learning under the largest possible set which includes all strategic opponents—i.e., seeks robustness to ‘everything’. This provides a novel utilitarian interpretation of the sense in which adversarial learning is worst-case. Based on this, our key modeling point is that prior knowledge regarding the opponent’s possible incentives can be used to reduce this large set. Since a smaller set means less to defend against, this has potential for reducing the unnecessary costs due to over-conservativeness. And since an empty set means no incentives and hence recovers the clean setting, strategic modeling enables a continuum of learning settings that interpolate between clean and adversarial, and thus provides the flexibility to extend beyond these two extremes. Our goal is to initiate the study of robust learning in these more nuanced cases.
@@
Towards this, we propose an approach for learning strategically-robust classifiers, which we refer to as strategic training. Our method builds on standard techniques from adversarial training and adapts them to handle strategic opponents with unknown utilities. Focusing on a large class of multi-targeted strategic opponents, we propose an implementation of such attacks that permit efficient training. The key challenge is that different uncertainty sets require different procedures; here we show that several natural families of strategic opponents admit simple and effective solutions. These include: (i) opponents who wish to preserve label semantics; (ii) opponents who benefit only from drastic label changes; and (iii) pre-specified targeted attacks. These are illustrated in Fig. 1, which shows how such strategic opponents lie between the adversarial and clean settings.
@@
To demonstrate the effectiveness of our approach, we perform a thorough empirical evaluation using multiple datasets and architectures and in various strategic settings. We begin by showing that adversarial training is suboptimal against strategic opponents, and analyze its weaknesses. We then show the merits of strategic training against known opponents, and study the risks of learning under misspecified incentives. Our main insight here is that the degree to which strategic training is useful depends on the relation between the structure of the opponent’s incentives and the structure of the learning task; for example, whether the opponent gains from semantic attacks (e.g., $\mathrm{dog}\mapsto\mathrm{cat}$) or from anti-semantic attacks (e.g., $\mathrm{dog}\mapsto\mathrm{truck}$). This motivates our analysis of learning under several natural strategic opponent classes, where despite their large uncertainty, the structure of incentives permits effective training. Results show that strategic training is much more effective against strategic opponents than adversarial training, while generally maintaining better clean accuracy. Code is available at https://github.com/maayango285/Adversaries-With-Incentives.
@@ 2 Related work
Adversarial examples are maliciously crafted inputs designed to make classifiers err. Amounting evidence suggests that imperceptible (but carefully chosen) noise suffices to cause missclasification [41, 15]. The process of constructing adversarial examples is termed an ‘adversarial attack’, and many such methods have been developed [15, 7, 10, 28]. Given the high susceptibility of neural networks to such attacks, there has been much interest in developing learning methods that provide adversarial robustness. One popular approach, and our focus herein, is adversarial training [15, 28], which modifies the training objective to include adversarial examples instead of raw ones. Adversarial training has been shown to be effective in improving adversarial accuracy; however, evidence suggests that a persistent trade off exists between robustness and accuracy on clean inputs [42]. Despite new insights [44] and recent advances [47, 43], this remains to be a major weakness.
@@
Our work draws inspiration from the fast-growing literature on strategic classification [17, 6, 16]. Works in this field have studied both theoretical [40, 46] and practical [26, 27] aspects of learning under strategic behavior, as well as broader notions such as transparency and information asymmetry [9, 14, 4, 3, 22], social implications [31, 26, 25, 48, 13], and causality [30, 8, 20, 29], among others (see [36]). Strategic classification studies learning in a setting where users can ‘game’ the system by modifying their features, at a cost, to obtain favorable predictions. Motivated by tasks such as loan approval, job hiring, college admissions, and welfare benefits, standard strategic classification considers binary tasks in which positive predictions (${\hat{y}}=1$) are globally better for users than negative (${\hat{y}}=0$). Hence, since all users wish to obtain ${\hat{y}}=1$, negative users ($y=0$) act ‘against’ the system, whereas positive users ($y=1$) are cooperative. Several works consider a more generalized setting in which all users can be adversarial-like [27, 40], but are still restricted to binary labels. To the best of our knowledge, our work is the first to consider strategic behavior in a multiclass setting, which enables us to establish a more deeper connection between strategic and adversarial learning.
@@
Many subfields within machine learning are concerned with providing worst-case robustness. These include robust statistics (e.g., [5]), distributionally robust optimization (e.g., [12]) and adversarial learning—all of which rely heavily on minimax formulations. In adversarial learning, some efforts have been made to promote the less restrictive notion of average-case robustness by replacing the max with an expectation over a prior [33, 34]. The idea to replace an adversarial opponent with a strategic (or rational) one has been considered to some extent in the domain of cryptographic proof systems [1]. Within learning, the only work we are familiar with which discusses a possible gap between worst-case and strategic guarantees is [32] who study no-regret learning in games. Our work proposes a strategic interpretation of worst-case robustness, achieved by modeling uncertainty as being over a set of strategic opponents. This draws a connection to the recent work of [35] who study minimax optimization in standard strategic classification under uncertainty in user costs.
@@ 3 Preliminaries
Let $x\in\mathbb{R}^{d}$ denote inputs and $y\in[K]$ denote categorical labels. We follow the conventional setup of supervised learning and assume input-label pairs $(x,y)$ are sampled iid from some unknown joint distribution $D$. Given a model class $F$, the standard goal in learning is to find some $f\in F$ that maximizes expected accuracy. For the 0-1 loss defined as $\ell(y,{\hat{y}})=\mathds{1}{\{{y\neq{\hat{y}}}\}}$, this amounts to solving:
@@
in which the attack $\delta^{\mathrm{adv}}$ is made explicit. This is the formulation that is typically optimized in practice (e.g., as in PGD [28]), often with replacing $\ell$ with a differentiable surrogate loss (e.g., cross-entropy). We refer to the expected error term in Eq. (1) as clean error and in Eq. (2) as adversarial error, and similarly define clean and adversarial accuracy.
@@ 3.1 Learning against strategic opponents
Thus, strategic opponents have the same capacity as adversarial opponents, but admit flexible objectives. Our goal in learning will be to defend against such attacks, a notion we term strategic robustness.
@@
Multi-targeted strategic attacks.
The strategic opponents we will focus on in this paper are based on the common notion of targeted attacks, but considered from a novel utilitarian perspective. [^1] Targeted attacks are typically employed post-hoc to demonstrate susceptibility to attacks against any arbitrary target. In contrast, we consider strategic opponents with a-priori preferences regarding the possible targets. Given a target class ${\bar{y}}\in[K]$, a targeted attack on $x$ is any $\delta\in\Delta$ that gives $f(x+\delta)={\bar{y}}$. Define a (multi)-targeted strategic opponent as one which gains utility from an attack only if it succeeds in flipping predictions to particular, label-dependent target(s). This captures settings such as our road signs example in which for each object of label $y$, only a subset of target classes ${\bar{y}}$ are beneficial. Formally, a targeted strategic opponent is one whose utility is $u(y,y^{\prime})=1$ if $y^{\prime}$ is a valid target for inputs with label $y$, and 0 otherwise (where again we assume $u_{yy}=0$). We refer to these as 0-1 utilities.
  • $k$-hot: An opponent is $k$-hot if each class $y$ has exactly $k$ viable targets; i.e., $\sum_{y^{\prime}}u(y,y^{\prime})=k\,\,\forall\,y$.
  • Semantic: If classes can be partitioned semantically (e.g., animals vs. vehicles, road sign types), an opponent is semantic if for all $y$, all viable targets are of the same semantic type (e.g., $\mathrm{dog}\mapsto\mathrm{cat}$).
  • Anti-semantic: Similarly, an opponent is anti-semantic if it benefits from changing label semantics, i.e., if for all $y$, all viable targets are of a different semantic type (e.g., $\mathrm{dog}\mapsto\mathrm{truck}$).
4 The drawbacks of adversarial training in strategic environments

### v3

@@ Abstract
Adversarial training aims to defend against adversaries: malicious opponents whose sole aim is to harm predictive performance in any way possible. This presents a rather harsh perspective, which we assert results in unnecessarily conservative training. As an alternative, we propose to model opponents as simply pursuing their own goals—rather than working directly against the classifier. Employing tools from strategic modeling, our approach enables knowledge or beliefs regarding the opponent’s possible incentives to be used as inductive bias for learning. Accordingly, our method of strategic training is designed to defend against all opponents within an ‘incentive uncertainty set’. This resorts to adversarial learning when the set is maximal, but offers potential gains when the set can be appropriately reduced. We conduct a series of experiments that show how even mild knowledge regarding the opponent’s incentives can be useful, and that the degree of potential gains depends on how these incentives relate to the structure of the learning task.
@@ 1 Introduction
The goal of adversarial learning is to train classifiers that are robust to adversarial attacks, defined as small input perturbations crafted to fool the classifier into making erroneous predictions (Szegedy et al., 2014; Goodfellow et al., 2015). As its name suggests, adversarial learning seeks to defend against an adversary—an ominous opponent whose actions are intended to induce maximal harm to the classifier’s performance. Learning in this setting can therefore be modeled as a zero-sum game between the learner and the adversary, in which each player’s gain translates to an equivalent loss for the other. This lends to formulating adversarial training as a minimax optimization problem in which the learner (min) controls model parameters, and the adversary (max) controls bounded-norm additive noise terms that are applied to inputs in response to the chosen parameters (Madry et al., 2018).
@@
Since learning seeks to obtain correct predictions, adversaries are modeled as gaining from any erroneous prediction—regardless of its type. This gives the adversary much flexibility: while there is typically only a single (or perhaps a few) correct labels, most labels are wrong, and each of these becomes a viable target. Thus, to obtain robustness, learning must defend simultaneously against all attacks targeting any incorrect label. In principle, this approach is sound—but its implementation requires placing demanding restrictions on the learning objective, whether explicitly or implicitly (Roth et al., 2020). These make robustness attainable, but at a cost; for example, it is well-known that adversarially trained models suffer from deteriorated generalization (Schmidt et al., 2018; Zhai et al., 2019) and reduced performance on clean (i.e., non-adversarial) data (Tsipras et al., 2019).
@@
In this paper, we argue that such costs can potentially be reduced by injecting into the learning objective a notion of the opponent’s incentives—i.e., what the opponent wants—which we propose to use as inductive bias. Notice that while conventional adversarial training considers the prospective harm of an attack, it lacks to consider the possible motives behind it: the fact that a panda can be turned into a gibbon does not provide grounds for why (nor whether) an opponent would wish to do so. Alternatively, our working assumption is that realistic opponents are much more likely to simply promote their own self-interests, rather than to always and purposefully counter those of learning—as is the working assumption of adversarial training. Thus, if we have some knowledge or beliefs regarding the opponent’s interests, even if quite general, then this can and should be exploited to improve prediction.
@@
Figure 1: Strategic opponents benefit only from certain types of attacks, such as preserving label categories (semantic), changing labels drastically (anti-semantic), admitting preferences (preference order), or mapping each class to a specific target (1-hot). As such, strategic learning offers a flexible middle ground between fully adversarial and clean settings. Our strategic training approach enables to encode prior knowledge regarding the opponent’s possible incentives into the learning objective.
Borrowing from the literature on strategic classification (Hardt et al., 2016; Brückner et al., 2012; Levanon & Rosenfeld, 2021), our approach is to model the opponent as a strategic agent whose actions are intended to maximize its own utility (rather than to generally minimize accuracy). This lends to our proposed formulation of learning as a non-zero sum game in which the goal is to defend against strategic attacks, i.e., those from which the opponent gains. The benefit of defending against strategic targets (rather than all targets) is that training becomes less constrained, and so robustness can be obtained with minimal sacrifice to clean accuracy. We propose a learning framework for training classifiers that are robust to strategic input manipulations, i.e., that attain strategic robustness.
The first step to strategic robustness is to ask: what determines the opponent’s utility? In some cases, the utility structure is apparent, such as when classes admit an ordering. Consider for example that sellers in online market platforms will likely try to make their products appear higher quality (and not lower) or more expensive (and not cheaper). Another example is that spoofing or impersonation attacks aim for higher clearance levels (and not lower). Other cases can be more nuanced; for example, an attacker who seeks to disrupt traffic by hampering with road signs will benefit more from some changes (e.g., ‘stop’ to ‘yield’) than from others (e.g., ‘no turn left’ to ‘right turn only’). Structure can also derive from the task at hand. For example, risk-averse opponents who fear detection may prefer milder attacks that preserve class semantics (e.g., one car model to another) since these are harder to detect. Contrarily, an opponent that aims to cause maximal harm may focus instead on anti-semantic attacks (e.g., car to toaster oven) if these inflict more damage. A final alternative is to infer utilities from past attack data, when available; we explore this idea initially in Appendix C.6.
Figure 2: Conjectured outcomes of strategic training (illustration).
Learning against strategic opponents therefore requires some knowledge or beliefs regarding possible incentives. Our proposed framework allows to encode these with an ‘incentive uncertainty set’, and learn classifiers that are robust to strategic attacks within this set. Our first observation is that when the uncertainy set is maximal, strategic and adversarial robustness coincide. Since an empty set reverts to standard (non-robust) training, strategic learning provides a means to interpolate between clean and adversarial performance in an informed manner. Smaller uncertainty sets enable higher clean accuracy, and improve robustness if they include the true incentives, but risk misspecification; larger sets reduce this risk, but at the cost of lower clean accuracy. By providing control over uncertainty, our goal is to allow the learner to make deliberate and calculated decisions about robustness in order to balance between potential gains and risks. This idea is illustrated in Fig. 2.
@@
Our approach builds on standard techniques from adversarial training and adapts them to handle strategic opponents with either known or uncertain utilities. Focusing on a large class of multi-targeted strategic opponents, we propose an implementation of such attacks that permits efficient training. The key challenge is that different uncertainty sets require different procedures; here we show that several natural families of strategic opponents admit simple and effective solutions. These include: (i) opponents who wish to preserve label semantics; (ii) opponents who benefit only from drastic label changes; (iii) opponents with a preference order over target classes; and (iv) pre-specified targeted attacks. These are illustrated in Fig. 1, which shows how such strategic opponents lie between the adversarial and clean settings. Our broader goal is to initiate the study of robust learning in these more nuanced cases, and motivate the exploration of others that lie between the two current well-known extremes.
@@
To demonstrate the effectiveness of our approach, we perform a thorough empirical evaluation using multiple datasets and architectures and in various strategic settings. We begin by showing that adversarial training is suboptimal against strategic opponents, and analyze its weaknesses. We then show the merits of strategic training against known opponents, and study the risks of learning under misspecified incentives. Our main insight here is that the degree to which strategic training is useful depends on the relation between the structure of the opponent’s incentives and the structure of the learning task. This motivates our analysis of learning under several natural strategic opponent classes. Results show that strategic training is much more effective than adversarial training against strategic opponents, while generally maintaining better clean accuracy. Results also show that the risks of learning under misspecified uncertainty sets exist, but are reasonable and often manageable. Together, these suggest that strategic modeling offers a valuable tool for training more precisely-defined robust classifiers. Code is available at https://github.com/maayango285/Adversaries-With-Incentives.
@@ 2 Related work
Adversarial examples are maliciously crafted inputs designed to make classifiers err. Amounting evidence suggests that imperceptible (but carefully chosen) noise suffices to cause misclassification (Szegedy et al., 2014; Goodfellow et al., 2015). The process of constructing adversarial examples is termed an ‘adversarial attack’, and many such methods have been developed (Goodfellow et al., 2015; Carlini & Wagner, 2017; Dong et al., 2018b; Madry et al., 2018). Given the high susceptibility of neural networks to such attacks, there has been much interest in developing learning methods that provide adversarial robustness. One popular approach, and our focus herein, is adversarial training (Goodfellow et al., 2015; Madry et al., 2018), which modifies the training objective to include adversarial examples instead of raw ones. Adversarial training has been shown to be effective in improving adversarial accuracy; however, evidence suggests that a persistent trade off exists between robustness and accuracy on clean inputs (Tsipras et al., 2019). Despite new insights (Yang et al., 2020) and recent advances (Zhang et al., 2019; Wang et al., 2019), this remains to be a major weakness.
@@
Our work draws inspiration from the fast-growing literature on strategic classification (Hardt et al., 2016; Brückner et al., 2012; Großhans et al., 2013). Works in this field have studied both theoretical (Sundaram et al., 2023; Zhang & Conitzer, 2021) and practical (Levanon & Rosenfeld, 2021; 2022) aspects of learning under strategic behavior, as well as broader notions such as transparency and information asymmetry (Dong et al., 2018a; Ghalme et al., 2021; Bechavod et al., 2022; Barsotti et al., 2022; Jagadeesan et al., 2021), social implications (Milli et al., 2019; Levanon & Rosenfeld, 2021; Lechner & Urner, 2022; Zhang et al., 2022; Estornell et al., 2023), and causality (Miller et al., 2020; Chen et al., 2023; Horowitz & Rosenfeld, 2023; Mendler-Dünner et al., 2022), among others (see Rosenfeld (2024)). Strategic classification studies learning in a setting where users can ‘game’ the system by modifying their features, at a cost, to obtain favorable predictions. Motivated by tasks such as loan approval, job hiring, college admissions, and welfare benefits, standard strategic classification considers binary tasks in which positive predictions (${\hat{y}}=1$) are globally better for users than negative (${\hat{y}}=0$). Hence, since all users wish to obtain ${\hat{y}}=1$, negative users ($y=0$) act ‘against’ the system, whereas positive users ($y=1$) are cooperative. Several works consider a more generalized setting in which all users can be adversarial-like (Levanon & Rosenfeld, 2022; Sundaram et al., 2023), but are restricted to binary labels. Most works in this space are theory-oriented, focus on linear classifiers, and support only continuous tabular features. To the best of our knowledge, our work is the first to consider strategic behavior in a multiclass setting over complex inputs, which enables us to establish a deeper connection between strategic and adversarial learning.
@@
Many subfields within machine learning are concerned with providing worst-case robustness. These include robust statistics (e.g., Bickel (1981)), distributionally robust optimization (e.g., Duchi & Namkoong (2021)) and adversarial learning—all of which rely heavily on minimax formulations. In adversarial learning, some efforts have been made to promote the less restrictive notion of average-case robustness by replacing the max with an expectation over a prior (Rice et al., 2021; Robey et al., 2022). The idea to replace an adversarial opponent with a strategic (or rational) one has been considered to some extent in the domain of cryptographic proof systems (Azar & Micali, 2012). Within learning, the only work we are familiar with which discusses a possible gap between worst-case and strategic guarantees is Piliouras et al. (2022) who study no-regret learning in games. Our work proposes a strategic interpretation of worst-case robustness, achieved by modeling uncertainty as being over a set of strategic opponents.
@@ 3 Preliminaries
Let $x\in\mathbb{R}^{d}$ denote inputs and $y\in[K]$ denote categorical labels. We follow the conventional setup of supervised learning and assume input-label pairs $(x,y)$ are sampled i.i.d. from some unknown joint distribution $D$. Given a model class $F$, the standard goal in learning is to find some $f\in F$ that maximizes expected accuracy. For the 0-1 loss defined as $\ell(y,{\hat{y}})=\mathds{1}{\{{y\neq{\hat{y}}}\}}$, this amounts to solving:
@@
in which the attack $\delta^{\mathrm{adv}}$ is made explicit. This is the formulation that is typically optimized in practice (e.g., as in PGD (Madry et al., 2018)), often by replacing $\ell$ with a differentiable surrogate loss (e.g., cross-entropy). We refer to the expected error term in Eq. (1) as clean error and in Eq. (2) as adversarial error, and similarly define clean and adversarial accuracy as one minus error.
@@ 3.1 Learning against strategic opponents
Thus, strategic opponents have the same capacity as adversarial opponents, but admit flexible objectives. In Sec. 5 we show how learning can defend against such attacks and attain strategic robustness.
Multi-targeted strategic attacks.
The strategic opponents we will focus on in this paper are based on the common notion of targeted attacks, but considered from a novel utilitarian perspective. [^1] Targeted attacks are typically employed post-hoc to demonstrate susceptibility to attacks against any arbitrary target. In contrast, we consider strategic opponents with a-priori preferences regarding the possible targets. Given a target class ${\bar{y}}\in[K]$, a targeted attack on $x$ is any $\delta\in\Delta$ that gives $f(x+\delta)={\bar{y}}$. Define a (multi)-targeted strategic opponent as one which gains utility from an attack only if it succeeds in flipping predictions to particular, label-dependent target(s). This captures settings such as our road signs example in which for each object of label $y$, only a subset of the classes $1,\dots,K$ are beneficial as targets. Formally, a targeted strategic opponent is one whose utility is $u(y,y^{\prime})=1$ if $y^{\prime}$ is a valid target for inputs with label $y$, and 0 otherwise (where again we assume $u_{yy}=0$). We refer to these as 0-1 utilities.
  • $k$-hot: An opponent is $k$-hot if each class $y$ has exactly $k$ viable targets; i.e., $\sum_{y^{\prime}}u(y,y^{\prime})=k\,\,\forall\,y$. These will serve us as a simple model for interpolating between clean and adversarial settings.
  • Semantic: In settings where classes admit a meaningful partitioning (e.g., animals vs. vehicles, road sign types), we say an opponent is semantic if for all $y$, all viable targets are of the same semantic type (e.g., $\mathrm{dog}\mapsto\mathrm{cat}$). Our empirical analysis reveals that adversarial attacks tend to preserve semantics; thus, semantic opponents typically act ‘as if’ they were adversarial—but are not.
  • Anti-semantic: In contrast, we say an opponent is anti-semantic if it benefits from changing label semantics, i.e., if for all $y$, all viable targets are of a different semantic type (e.g., $\mathrm{dog}\mapsto\mathrm{truck}$). We use these to move away from conventional adversaries and towards more lenient opponent types.
  • Preferences: Finally, an opponent adheres to a preference order if there is some (partial) ordering over classes such that $y\preceq y^{\prime}$ means $y^{\prime}$ is preferable to $y$, i.e., $u(y,y^{\prime})=\mathds{1}{\{{y\preceq y^{\prime}}\}}$. These are useful for settings in which classes are associated with price, quality, or popularity, which in turn effect the opponent’s gains, such as in recommender systems or online market platforms.
@@
4 Drawbacks of adversarial training in strategic environments